Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SCS SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
Technical integration and coordination SASE1 Status and SASE3 Readiness
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 592     Entry time: 03 Mar 2025, 09:48
Author: Frederik Wolff-Fabris 
Location: General 
Severity: Information 
Subject: Notes XO Integration Meeting 03.03.2025 
General Announcement:
- For next Friday 07.03 UOM a jitter report from instruments is expected together with presentations from groups using beam during this FD week.

CTRL:
- Nothing to report.

DA:
- Not present.
 
EEE:
- Nothing to report.

VAC:
- Issue with vacuum at HED Mono on Sunday morning 06:40AM: spike on vacuum happened (needs investigation with HED and EEE to further understand it). Recovered half way by HED colleagues. 
Please do not hesitate in calling VAC-OCD in such situations. The V0 closed and was masked at MPS - now it is back unmasked (Desy-MXL will investigate).

XPD:
- Nothing to report.

DET:
- Nothing to report.

ITDM: 
- Nothing to report.

XRO:
- Nothing to report.

LAS:
- Not present.

UNSYS:
- Nothing to be reported.

SPB:
- Energy drop on Saturday 02:20AM (bunches not lasing) was seen (fixed); vacuum spike in the chamber being investigated by SPB/SFX.

FXE:
- Nothing to report.

MID:
- Nothing to report.

HED:
- Apart of the VAC, all fine.

SCS:
- Discussion with BKR will be done.

SQS:
- Nothing to report.

SXP:
- Nothing to report.
ELOG V3.1.4-7c3fd00